OpenStudy (anonymous):

Please help i don't understand how to do this ill give medal for best answer :) 1.Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qmetal my data Measured mass metal (Zinc) 41.664g Distilled water measurement 26mL Distilled water temperature 25oC Temperature of metal 100.8oC Temperature of mixture 34.8oC

OpenStudy (frostbite):

Okay weird question they want enthalpy in the question, but just to please your teacher(s) we can integrate enthalpy into our equations: First law of thermodynamics: \[\Large \Delta U = q+w\] Lets assume constant volume and no additional work happens: \[\Large \Delta U=q\] Lets assume constant pressure too: \[\Large \Delta H = \Delta U = q\] We assume that the only think working on the surroundings is the heat transfer from the system, so we can write: \[\Large \Delta H=\Delta U=q=c \times m \times \Delta T\] Use the equation above and calculate the, internal energy chance / enthalpy / heat process.

OpenStudy (frostbite):

that the only thing working*